please help it say use the substitution method to solve the linear system
x+2y=-5
4x-3y=2

32-2y=4
x+ 3y=5

3x+y= -2
x+3y=2

Answer by checkley71(8403)

X=-5-2Y THEN 4(-5-2Y)=2 OR -20-8Y=2 OR -8Y=22 OR Y=-22/8 OR Y=-11/4
X=-5-2(11/4) OR X=-5-22/4 OR X=(-20-22)/4 OR X=-42/4 OR X=-21/2